OpenStudy (adamconner):

what is the integral of [csc(sinx)]^2 cosx dx, when u=sinx

hartnn (hartnn):

what did u get after plugging in u= sin x ?? du =.. ?

OpenStudy (adamconner):

cosx

hartnn (hartnn):

du = cos x dx to be precise so, whats your integral now ?

hartnn (hartnn):

in terms of u only

OpenStudy (adamconner):

[cscx(u)]^2 du..

hartnn (hartnn):

correct! can you find that integral ? its standard

OpenStudy (adamconner):

do i break up the ^2 for the u and csc x?

hartnn (hartnn):

do you know the derivative of cot x ?

OpenStudy (adamconner):

csc^2 x

hartnn (hartnn):

its - csc^2 x right ?? \((d/dx)\cot x = -\csc^2x \\ \implies \int \csc^2xdx = -\cot x +c\) got this ?

OpenStudy (adamconner):

yea i got it.

hartnn (hartnn):

thats it! final answer is -cot x +c :)

hartnn (hartnn):

now only thing remain is to plug back in for 'u'

OpenStudy (adamconner):

-cot(sinx) +c

hartnn (hartnn):

oh, typo, so you got -cot u+c right ??

hartnn (hartnn):

and u = sin x gives you -cot (sin x) + c

hartnn (hartnn):

yes :)

OpenStudy (adamconner):

yea thats what i got. Thanks!

hartnn (hartnn):

welcome ^_^

OpenStudy (adamconner):

one question. why didnt we square the sin x. we only did the csc x?? @hartnn

hartnn (hartnn):

we plugged in u = sin x so we had csc^2 u du where was the need to square the sin x ?

OpenStudy (adamconner):

but isnt the square for both csc(sinx) so we distribute the ^2 to both terms right? so we plugged in u for sin. so why not csc^2 u^2?

hartnn (hartnn):

oh, its not PRODUCT of function ! not like (ab)^2 = a^2 b^2

hartnn (hartnn):

its composite function, csc (sin x) ------> like f(g(x))

hartnn (hartnn):

and [f(g(x))]^2 = f^2 (g(x))

OpenStudy (adamconner):

ooh. now it makes more sense! thanks again :)

hartnn (hartnn):

welcome ^_^